# Learning from Interaction: A Survey

### v1

@@ 5.5 Robustness Scaling
Feedback robustness scaling requires the environment to provide more stable and reliable reward signals. Regarding the robustness of the reward itself, signals may be noisy, or face the challenge of reward hacking, which refers to agents learning undesirable and tricky behavioral patterns that obtain high reward without achieving the intended goals (miao2024inform; liu2024rrm; farquhar2025monamyopicoptimizationnonmyopic; fu2025reward; tarek2025reward). For the former issue, some approaches generate soft probabilistic rewards with generative verifiers to mitigate noise in rewards (lin2024navigating; su2025crossingrewardbridgeexpanding). And as for the latter, frameworks such as MONA (farquhar2025monamyopicoptimizationnonmyopic) try to mitigate reward hacking problem by evaluating the future utility of actions through an overseer, thus constraining unstable behaviors while preserving explainability. huang2025reinforcement develops a reward hacking defense rubric that penalizes sycophantic praise towards user prompts and overly flattering self-assessments in responses, encouraging the model to produce more substantive content.
Apart from reward robustness, interaction robustness at environment-level is also crucial, since it is common for environmental instabilities (e.g., delays, crashes, corrupted tool outputs) to undermine feedback reliability and degrade the training process. To handle such failures, Trinity-RFT (pan2025trinityrftgeneralpurposeunifiedframework) proposes asynchronous inference and retry mechanisms, while tongyidr employs caching, retrying failed calls, and switching to similar providers to prevent corrupted trajectories. Looking ahead, future reward design should prioritize both efficacy and the prevention of hacking through tricky patterns, and greater attention should be paid to the design of more robust system.
6 Implementation Frameworks
There exist diverse implementations of simulated environments that vary in modality and complexity.
Visual Environments
For 2D scenarios, tasks in grid-based environments can be symbolic or pixel-level, often associated with game-playing (chevalier2023minigrid; bellemare2013arcade). Works such as WebArena (zhou2023webarena), Mind2Web (deng2023mind2web), and VisualWebArena (koh2024visualwebarena) extend agent capabilities to more realistic, web-based environments. ARE (andrews2025arescalingagentenvironments) further builds upon this line by introducing a mobile-style setting that integrates multiple apps and tools, incorporating time control and a structured verifier to support scalable evaluation. In 3D scenarios, environments emphasize more about embodied interaction, realistic physics and visual perception. More general simulators such as Habitat (savva2019habitat) and ThreeDWorld (gan2020threedworld) support more actions such as navigation and manipulation, while domain-specific worlds develop the specific aspect. For example, MineRL (guss2019minerl) and MineDojo (fan2022minedojo) exploit the flexibility of Minecraft. Besides, Household benchmarks such as ALFRED (shridhar2020alfred) and EmbodiedQA (das2018embodied) further evaluate the grounding function of language in photo-realistic 3D spaces, requiring agents to follow instructions, answer questions, or perform multi-step tasks.
Text-based Environments
Text-based environments emphasize more on reasoning and decision-making from natural language descriptions. TextWorld (cote2018textworld) proposes a framework to generate the text games with different goal-driven challenges. Reddit-RL-simulator (he-etal-2016-deep-reinforcement; chan-king-2018-thread) implements an RL environment for iteratively tracking and recommending popular discussion threads on Reddit. ALFWorld (shridhar2020alfworld) adapts ALFRED tasks into textual form, and JerichoAgentBench (liu2023agentbench) extends the Jericho suite (hausknecht2020interactive) with annotated benchmarks and concrete objectives. Such text environments stress the agent’s ability to infer latent dynamics under partial observability, making them a natural testbed for language-based decision-making.
@@ 8 Conclusion
The era of experience makes environments central to the development of LLM agents, positioning them as active producers of experiential data and underscoring the growing need for scaling environments to create a more complex, realistic, and richly interactive world. From a pioneering environment-centric perspective, our survey proposes a unified taxonomy that organizes representative work across the GEF loop (task generation, task execution, and feedback), together with evaluation, implementation, and applications. Besides, we surface key challenges for advancing agent intelligence, including the asymmetry between generators and verifiers and the construction of more open, large-scale multi-agent environments, thereby providing insights for future research on agentic systems.
@@ Acknowledgments
Appendix A Conceptual Framework
@@
Appendix B Evaluation Benchmarks
@@
Appendix C Key Applications
@@
Recent progress in agentic systems, typically built on state-of-the-art LLM families such as GPT (openai2025introducing), Claude (anthropic2025claude4), Gemini (team2024gemini), LLaMA (touvron2023llama), and Qwen (bai2023qwen), is increasingly driven by interactions with dynamic and multifaceted environments.
Tool-use Environments
@@
Tool-use environments expose APIs and function calls as structured action spaces, and many LLMs now natively support function invocation, thereby extending reasoning with external tools (OpenAIFunctionCalling2023; anthropic2024mcp; wu2025webwalkerbenchmarkingllmsweb; mastouri2025makingrestapisagentready; luo2025mcp; wang2025mcp; fan2025mcptoolbench++).
Coding Environments
@@
Coding environments leverage repositories, test frameworks, and IDE integration to support long-horizon programming. Within these settings, systems such as Qwen3-Coder (qwen3coderblog2024) and Claude 4 (anthropic2025claude4) demonstrate reliability in code editing and debugging. ReTool (feng2025retoolreinforcementlearningstrategic) further integrates code-interpreter execution into the reasoning loop, enabling agents to exhibit code self-correction and adaptive tool selection.
GUI Environments
@@
Web navigation (browser control) environments build on HTML/DOM structures to support tasks such as browsing, form filling, and transactions (lai2024autowebglm; AgenticSeek2024; AutoGPT2023). GUI environments extend these to graphical user interfaces on desktops and mobile devices (qin2025uitars; wang2025opencua; ye2025mobileagentv3; lu2025guiodyssey; liu2024autoglm; sun2022metagui; huang2025guikvefficientguiagents). Beyond these, more comprehensive platforms (wang2025uitars2; xie2024osworld; OpenAICUA2025) integrate terminals, operating systems, applications, and APIs to create more interactive, open-ended, and realistic environments, fostering the development of more advanced computer-use agents (CUAs).
Deep Research Environments
To nurture the next generation of more powerful research agents, deep research environments demand stronger long-context reasoning and more robust retrieval capabilities. Systems such as Gemini 1.5 Pro (google2024geminideepresearch) and OpenAI’s deep research agents (openai2025deepresearch) demonstrate that extended context windows enable sustained, in-depth analysis, while effective retrieval pipelines help mitigate distraction and context dilution.

### v3

@@ 4.5 Robustness Scaling
Feedback robustness requires the environment to provide stable and reliable reward signals. Reward-level robustness concerns the stability of the reward itself, since such signals may be inherently noisy or susceptible to hacking (agents learning undesirable and tricky behavioral patterns that obtain high rewards without achieving the intended goals). To mitigate noise, lin2024navigating and su2025crossingrewardbridgeexpanding generate soft probabilistic rewards with generative verifiers. To prevent reward hacking, farquhar2025monamyopicoptimizationnonmyopic evaluates the future utility of actions through an overseer, constraining unstable behaviors while preserving explainability. huang2025reinforcement develops a defense rubric that penalizes sycophantic praise toward user prompts and overly flattering self-assessments in responses, encouraging the model to produce more substantive content. Environment-level robustness is also crucial for agents, as environmental instabilities (e.g., delays, crashes, corrupted tool outputs) may also undermine feedback reliability and degrade the training process. To handle such failures, Trinity-RFT (pan2025trinityrftgeneralpurposeunifiedframework) proposes asynchronous inference and retry mechanisms, while tongyidr employs caching, retrying failed calls, and switching to similar providers to prevent corrupted trajectories. Both approaches contribute underlying frameworks that ensure stable feedback.
[Uncaptioned image] Future Frontiers
Future reward designs should prioritize both efficacy and the prevention of hacking through tricky patterns, with greater attention also given to the development of more robust feedback mechanisms and underlying frameworks.
5 Implementation
5.1 Environmental Designs
In this section, we discuss environmental design choices across domains. In Tool-Use scenarios, environmental designs prioritize expanding the interactivity width and depth. This involves supporting deeper tool call sequences to solve more complex tasks and incorporating more general, versatile tools to handle tasks across a broader range of domains.
Similarly, Deep Research scenarios focus on enabling agents to orchestrate diverse search and browsing APIs during task generation, as well as robust retrieval and reasoning through long contexts. For example, DR Tulu agent infrastructure (shao2025drtulu) optimizes a unified MCP-based tool backend through global caching and asynchronous process lock, supporting high-concurrency operations. ARE (andrews2025arescalingagentenvironments) further transforms tool-use benchmarks into separate applications, integrating mobile use and tool use into a unified environment.
In Software Engineering (Coding) scenarios, designs emphasize whether the agent can actually execute code and whether the process can be reproduced. Docker (merkel2014docker) is commonly utilized as the basis of execution sandbox, as it ensures consistency across different machines through standard Dockerfiles while also enabling the safe execution of code within an isolated virtual environment, preventing potential safety issues (yang2023intercodestandardizingbenchmarkinginteractive; jimenez2023swe; yang2025swesmithscalingdatasoftware).
Web Navigation scenarios also focus on realism and interactivity scaling. WebArena (zhou2023webarena) delivers websites in self-contained Docker images as well, avoiding unexpected content or configuration changes. WebShop (yao2023webshopscalablerealworldweb) offers a diverse range of semantic action space (searching queries and choosing buttons), with its modular architecture facilitating scaling to new web tasks and domains. VisualWebArena (koh2024visualwebarena) further introduces a novel Classifieds website built on OSClass, a robust open-source Content Management System (CMS), supporting visually grounded tasks such as posting, searching, and commenting, thereby offering a more interactive environment for user engagement.
In Computer Use scenarios, in addition to interactivity and realism, data collection efficiency and diversity integration (e.g., cross-OS tasks, heterogeneous tools) are also key concerns. OpenCUA (wang2025opencua) and OSWorld (xie2024osworld) provide infrastructures that support efficient task annotation and the real-time capture of user interactions. UI-TARS-2 (wang2025uitars2) unifies GUI operations and SDK function calls into a versatile platform, supporting a variety of task kinds and maintaining long-lived states across intricate, multi-step interactions.
In Embodied scenarios, tasks involve requiring agents to follow instructions, ground objects in photo-realistic 3D spaces, answer questions, or perform multi-step tasks (weihs2021visualroomrearrangement; das2018embodied). These tasks require environments to maintain real-world realism, such as actual grounding and collision avoidance. Frameworks like ALFWorld (shridhar2020alfworld) provide the necessary infrastructure for abstract reasoning and concrete execution. More recently, Genie 3 (parker2025genie3) has further improved real-time interactivity and long-horizon consistency, providing a fertile foundation for the advancement of future LLM agents.
5.2 Implementation Challenges
A key challenge in environmental design is the Generator-Verifier Asymmetry (wei2025asymmetry), which refers to the mismatch between the intelligence required for the generator and that required for the verifier. In easy-to-verify domains (e.g., mathematics and programming (wei2025browsecomp; jimenez2023swe; phan2025humanity)), task generation can be difficult (scaling up high-quality, non-trivial tasks is challenging), while verification is relatively easy and computationally inexpensive. In contrast, in easy-to-generate domains (e.g., creative writing and healthcare (lin2024wildbench; arora2025healthbench)), it is easy to propose open-ended and challenging tasks, but verification may involve subjectivity or require expert knowledge. Future environmental designs may allow the generator and verifier to co-evolve, tapping into the potential of using a powerful generator as a verifier to stimulate the development of even more powerful generators (huang2025r; hong2025cooper; lu2025urpo; chen2025self; wang2025socraticzerobootstrappingreasoning). See more details are in Appendix A.
@@ 7 Conclusion
The era of experience makes environments central to the development of LLM agents, positioning them as active producers of experiential data and underscoring the growing need for scaling environments to create a more complex, realistic, and richly interactive world. From a pioneering environment-centric perspective, our survey proposes a unified taxonomy that organizes representative work across the three-stage GEF loop (task generation, task execution, and feedback), together with implementation frameworks and key applications. Additionally, we surface key challenges and highlight future frontiers for advancing agent intelligence along each scaling dimension, offering valuable insights for future research on agentic systems.
@@ Acknowledgments
Appendix A Background
A.1 Scaling Laws for LLM Agents
A.2 Generator-Verifier Asymmetry
Figure 5: Generator-verifier asymmetry challenge.
A fundamental characteristic in many real-world tasks is the inherent Generator-Verifier Asymmetry (wei2025asymmetry), namely the mismatch between the intelligence required for generator, which generates (§2) or executes (§3) tasks, and that required for verifier, which provides feedback (§4). These two kinds of intelligence naturally form two axes critical to next-generation Agentic AI, as illustrated in Figure 5. From this perspective, scaling up environments essentially corresponds to scaling intelligence along the $x$-axis and the $y$-axis. Current progress in RL largely exploits the regime on the easy-to-verify side of this asymmetry. These Easy-to-Verify, Hard-to-Generate Domains include fields such as mathematics and programming (wei2025browsecomp; jimenez2023swe; phan2025humanity). For these domains, generating and solving a continual stream of high-quality, non-trivial tasks is challenging. In contrast, verification is objective and computationally inexpensive (e.g., via unit tests or exact match on mathematical results). This enables weak-to-strong supervision, where a simple verifier can provide accurate feedback to train a much stronger agent for solving hard tasks. On the contrary, the Hard-to-Verify, Easy-to-Generate Domains include areas such as creative writing, policy-making, or healthcare (lin2024wildbench; arora2025healthbench). For these easy-to-propose, open-ended tasks, verification is subjective, requires substantial expert judgment, or unfolds over long horizons, making high-quality feedback scarce and expensive. This bottleneck, corresponding to the upper-left region of coordinate system, poses more difficulty in modeling the environment, and rendering environment scaling more challenging yet offering greater potential for advancing agent capabilities. Notably, the asymmetry also presents an opportunity: if the generator’s stronger intelligence can be systematically leveraged to strengthen the verifier, so that it can supervise an even stronger generator, then such asymmetric property can be exploited to drive agents’ self-evolution (huang2025r; hong2025cooper; lu2025urpo; chen2025self; wang2025socraticzerobootstrappingreasoning).
A.3 Conceptual Framework
@@
Appendix B Comparative Analysis
A comparative analysis of methods for scaling task complexity and task diversity is presented in Tables 1 and 2.
Table 1: Comparative Analysis of Task Complexity and Performance Across Methods
Method Skill Dimension Task Structure Task Complexity Multi-Turn Performance
Width Depth
ToolBench (qin2023toolllm) Tool Use Sequential 1 - $\times$ -
$\tau$-bench (yao2024tau) Tool Use Sequential 1 -, 30 $\checkmark$ -
BFCL V3 (patil2025bfcl) Tool Use Sequential 1 -, 20 $\checkmark$ -
[…]
Notes: Width denotes the number of requirements that must be satisfied within a single task. Depth denotes the number of tool calls/steps per task, where the first value indicates the range of steps taken by the majority of tasks and the second value gives the maximum steps. Performance reports GAIA (mialon2023gaia) scores (if available). Multi-Turn indicates whether interaction between the user and the agent is multi-turn.
Table 2: Comparative Analysis of Environment/Task Diversity across Methods
Method Modality Skill Dimension # of Envs # of Tasks # of Trajs Source Performance
SWE ALF
SWE-Gym-32B (pan2025trainingsoftwareengineeringagents) Text Code; - 2,438 491 Real 20.6 -
R2E-Gym-32B (jain2025r2egymproceduralenvironmentshybrid) Text Code; - 8,135 3,321 Synth 34.4 -
SWE-Smith-32B (yang2025swesmithscalingdatasoftware) Text Code; - 50,137 5,016 Both 40.2 -
AgentGen-70B (hu2025agentgen) Text Planning; 592 11,840 7,246 Synth - 19.4
AgentGym-70B (xi2024agentgym)
Text +
Visual
Web; Embodied;
Code; Tool Use;
14 89 6,130 Both - 88.0
AgentBank-13B (song2024agentbank)
Text +
Visual
Web; Embodied;
Code; Tool Use;
16 - 51,287 Both - 72.4
Notes: SWE refers to SWE-bench (jimenez2023swe). ALF refers to ALFWorld (shridhar2020alfworld). Performance on out-of-domain tasks.
Appendix C Evaluation Benchmarks
@@
Appendix D Applications Across Domains
@@
Driven by interactions with dynamic and diverse environments, recent agentic systems built on state-of-the-art LLM families such as GPT (openai2025introducing), Claude (anthropic2025claude4), Gemini (team2024gemini), and Qwen (bai2023qwen) have become increasingly powerful and widely applicable.
Tool Use
@@
In tool-use scenarios, environments expose APIs and function calls as structured action spaces for agents, enabling them to reason with external tools through function calling or code generation (OpenAIFunctionCalling2023; anthropic2024mcp; wu2025webwalkerbenchmarkingllmsweb; mastouri2025makingrestapisagentready; luo2025mcp; wang2025mcp; fan2025mcptoolbench++).
Software Engineering
@@
In code generation or software engineering scenarios, environments leverage repositories, test frameworks, and IDE integration to cultivate agents’ long-horizon programming capabilities. Systems such as Qwen3-Coder (qwen3coderblog2024) and Claude 4 (anthropic2025claude4) demonstrate reliability in code editing and debugging. ReTool (feng2025retoolreinforcementlearningstrategic) novelly integrates code-interpreter execution into the reasoning loop, enabling agents to exhibit code self-correction and adaptive tool selection. SWE-smith (yang2025swesmithscalingdatasoftware) and R2E-Gym (jain2025r2egymproceduralenvironmentshybrid) develop scalable, executable code environments based on GitHub repositories, enabling the further evolution of SWE agents.
Web Navigation
@@
In web navigation scenarios, environments are typically built on HTML/DOM structures to support tasks such as browsing, form filling, and transactions (lai2024autowebglm; AgenticSeek2024; AutoGPT2023), requiring agents to click buttons and navigate between web pages. Graphical user interface (GUI) environments extend these tasks to desktops and mobile devices, further complicating task complexity (qin2025uitars; ye2025mobileagentv3; lu2025guiodyssey; liu2024autoglm; sun2022metagui; huang2025guikvefficientguiagents).
Computer Use
In more recent computer-use scenarios, environments integrate terminals, different operating systems, various applications, and APIs, making tasks much more interactive, open-ended, and realistic (wang2025opencua; wang2025uitars2; xie2024osworld; OpenAICUA2025; jia2025osworldmcp; xie2025scalingcomputerusegroundinguser).
Deep Research
In deep research scenarios, environments demand stronger long-context reasoning and more robust retrieval capabilities, as demonstrated by prominent long-form tasks like AstaBench (bragg2025astabench), DeepResearchBench (du2025deepresearchbench), and ResearchQA (yifei2025researchqa). By learning from the interactive experiences collected from these environments, agentic systems such as Tongyi DR (tongyidr), Gemini DR (team2024gemini), OpenAI DR (openai2025deepresearch), and DR Tulu!(shao2025drtulu) are able to conduct sustained, in-depth analysis with mitigated context dilution.
Appendix E Ethics and Safety Implications
Agent systems require robust ethical safeguards, interoperable standards, and alignment with emerging regulatory frameworks to enable responsible integration into the global technology ecosystem (howell2024policy). Some prior work focuses specifically on safety, such as mitigating inherent biases that produce discriminatory outputs or defending against malicious inputs that circumvent safety-alignment mechanisms (zhao2023art). However, new challenges are likely to emerge as agents accumulate experience through interactions with their environments and with other agents, and as their capabilities for autonomous task planning, external data access, and tool use continue to grow. For example, prompt injection or tool injection attacks on a single agent may lead to privacy breaches or system-wide contamination, while agents empowered to autonomously operate on external data can, in the absence of strict safeguards, cause irreversible damage by initiating fund transfers or deleting critical files (tang2025risks). Therefore, when developing scaled environments for future LLM-based agents, safety considerations must be treated as a central and indispensable element of system design.